OpenStudy (mhchen):

Level: HARD + Super medal On average, a noob gets 1 kill per 800 games. He plays 200 games in a day. What is the probability that the noob gets 2 or more kills per day?

OpenStudy (mhchen):

ideas: This is a poisson distribution

OpenStudy (mhchen):

I just solved it. 1 - poissCdf(0.25,0,1) Where 0.25 is 200/800

OpenStudy (mhchen):

anyone like medal?

OpenStudy (dumbcow):

ok using the poisson distribution we know that n = 200 \[\lambda = 200* \frac{1}{800} = \frac{1}{4}\] So the expected num of kills per day is 1/4 \[P(X = k) = \frac{e^{-\lambda} \lambda^k}{k!}\] To find probability of more than 2, we need to find prob of 0 kills and 1 kill \[P(X \ge 2) = 1 - [P(X=0) + P(X=1)]\]

OpenStudy (salty):

\(\lambda =200/800=1/4=0.25\) and yeah then we subtract probability of getting 0 and 1 from 1 so \(P(0)=\large\frac{(0.25)^0 \times e^{-0.25}}{0!}\) \(P(1)=\large\frac{(0.25)^1 \times e^{-0.25}}{1!}\) P(needed)=1-0.7788007 -0.1947 =0.0264

OpenStudy (mhchen):

Geez :b I underestimated you guys.

OpenStudy (salty):

just learned this stuff today thanks!

OpenStudy (mhchen):

Nice! we must be in the same class then.

OpenStudy (salty):

We started probability stuff 3 lectures ago and next up is conics